OpenStudy (anonymous):

find the equation of the line that is parallel to x=5y-4 and it passes through (3,-5) answer in standard form

OpenStudy (anonymous):

x-5y=18

OpenStudy (kainui):

What makes a line parallel to another line? They must have the same slope. So to help yourself out, put your equation into standard form, y=mx+b You know that m is your slope and b is your y-intercept. Does this help?

OpenStudy (anonymous):

Its slope

OpenStudy (anonymous):

Use the "point slope formula". Y1-y=m(X1-x)

OpenStudy (anonymous):

thank u